OpenStudy (anonymous):

how to solve 2 2/3 +1 3/8=

OpenStudy (nirmalnema):

convert it to improper fraction then add them.. 8/3 + 11/8 take LCM of 8and 3 (64 + 33 ) / 24 97/24 =covert it to mixed fraction 4 1/24

OpenStudy (ranga):

You can also add the whole number portion separately and the fraction part separately and then combine them in the end: 2 + 1 = 3 2/3 + 3/8 = 16/24 + 9/24 = 25/24 = 1 & 1/24 add them all: 4 & 1/24